OpenStudy (anonymous):

-3/4x = -5/6 Solve using the multiplication Method. (check your work)

myininaya (myininaya):

\[\frac{-3x}{4}=\frac{-5}{6}\] right? \[-3x*6=-5*4\] -18x=-20 x=-20/(-18) x=20/18 x=10/9

OpenStudy (anonymous):

the x is after the 3/4 ... :)

myininaya (myininaya):

\[\frac{-3}{4}*x=\frac{-3x}{4}\] samething

OpenStudy (anonymous):

okay.. hell I am sooo bad at math just had to make sure :) thank you !

myininaya (myininaya):

\[\frac{a}{b}*\frac{c}{d}=\frac{a*c}{b*d}\]